OpenStudy (anonymous):

For which equation would you subtract 3, and then multiply by 9 to solve? nine times k plus three equals forty the quantity k over nine all plus three equals forty nine times k minus three equals forty the quantity k over nine all minus three equals forty

OpenStudy (anonymous):

B

OpenStudy (anonymous):

it is because we follow the exact opp procedure when we try to find the equation

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

You want to get "k" by itself. Here's the equations written out: 9*k+3=40 k/9+3=40 9*k-3=40 k/9-3=40 Which one do you think?